OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = 13 and a13 = -59? -185 -179 -173 -167

OpenStudy (callisto):

common difference d = (-59-13)/ (13-1) = -72 / 12 = -6 32nd term = a1 + (n-1)d = 13 + (32-1) (-6) = 13 + (31)(-6) =???